A Complete Bibliography of Publications in Synthese, 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

1956 [105].

accuracy [72]. Acknowledgments [54]. action [94]. Activités [9, 70]. Activities [9, 70]. activity [78]. address [15, 57]. aesthetics [37]. After [56]. Albert [46, 45]. Ambiguité [125]. Ambiguity [125]. among [102]. analyse [102]. analysis [31, 36, 6, 102, 120]. application [12, 65, 21]. arithmetic [38]. Aspects [99, 90, 85, 20, 17]. Association [149]. Ausbildung [75]. Ausdruck [73]. authority [92]. automatic [129]. aux [59, 22]. Axiomatic [58]. axiomatique [58].

B [98]. background [89]. backgrounds [136]. be [154]. Bedeutung [128]. Begriffe [73]. Begriffsbildung [62]. behaviour [141]. Bemerkungen [51]. between [37]. Bibliography [106]. Binaire [125]. Binary [125]. biological [27]. biology [25, 154]. Book [41, 121, 156, 53, 13, 8]. Bourgin [116].

C [10]. Calcul [125]. Calculation [125]. Can [154]. capacity [26]. Causalité [130]. Causality [130]. chance [59]. chez [96, 102]. child [96]. chronogeometry [46]. classification [140]. Clay [42, 43]. cochlear [26]. Code [125]. Combination [68]. communication [27, 20, 17, 126, 22]. communicative [30]. Compared [110]. Comparée [110]. Concealment [34]. concept [12]. conceptions [112]. Concepts [73, 20, 21, 113]. Conceptualisation [62]. conditionals [6]. Congresses [138, 148]. Constructive [60]. constructives [60]. content [31]. contribution [22, 28]. Corrigendum [14]. counterfactual [6]. Criteria [36]. cultural [45]. Cybernetic [158]. Cybernetics [110, 32]. Cybernétique [110].

dans [65, 135, 131, 4]. Dantzig [150, 152, 151]. David [150, 152, 151]. day [79]. demand [142]. Denken [51]. Déscription [131, 131]. Descriptions [61]. Development [4, 81, 88, 93, 75, 100]. développement [4]. d’expression [83]. Didactic [69, 80]. Diesel [51]. Dieselben [51]. dimensions [72]. Dimensionsbegriffen [72]. d’Interprétation [135, 102]. dissolving [37]. Distances [118]. distinction [37]. Droit [135, 65]. Dutch [149].

ear [26]. Einfluss [75]. Einheitswissenschaft [72]. Einstein [46, 44, 45]. Elementarmathematik [109]. elements [88, 95, 102, 102]. emotional [96]. émotionnelles [96]. empirical [119]. Empiricism [4]. empirio [120]. empirio-semantical [120]. Empirisme [4]. endechy [21]. enfant [96]. enquête [65]. enterprise [120]. Entwicklung [81, 75]. equivalence [114]. especially [96, 102]. essence [145]. ethnologie [83]. ethnology [83]. Evidence [144]. exact [154]. Exacte [149]. Exaktheit [72]. excitation [25]. exemplification [3]. experimental [99]. Expression [73, 83]. Expressions [9, 70].

F [10]. familiar [49]. Final [35]. flax [52]. Formal [78, 94]. Formalised [143]. Formalisierung [72]. Formalization [72]. formation [75]. foundation [89]. Foundations [29]. French [40, 104, 64, 84, 61, 110, 60, 91, 65, 135, 59, 96, 83, 125, 131, 130, 87, 9, 70, 58, 4, 63, 102, 22, 132]. functions [81]. fundamental [20]. Funktionen [81]. future [51].

G. [10, 112, 153]. Gegenstandsbeziehing [128]. genèse [59]. genesis [59]. géometrie [40]. geometry [40]. Georges [116]. German [49, 73, 67, 81, 50, 48, 74, 62, 72, 128, 51, 75]. Gerrit [106, 103, 105, 104]. Gino [51, 51]. God [48]. Gottesbeweis [48]. grammar [66]. Griss [10]. growth [76].

Hans [2]. hasard [59]. head [24]. heuristic [76]. Heuristics [67]. Heuristik [67]. historical [64]. historique [64]. Hommage [104]. Hospinianus [50, 50]. Human [129, 20, 26]. hypotheses [39].

Ib [102]. Idea [59]. Idée [59]. ideological [89]. ihr [73, 75]. illustration [144]. im [128]. impact [107]. Inaugural [15]. induction [142]. influence [75]. informal [16]. Information [31, 131, 24, 124, 18, 21, 34, 26, 16, 28, 19, 131]. Informative [127]. intentional [114]. International [158]. Interpretation [135, 130, 126, 131, 130, 134, 92, 133, 131, 102, 129, 80, 124]. Interprétations [132, 134, 132]. interventions [87]. Introduction [123, 115, 23, 55, 19]. Introductory [57]. intuitionistic [10]. inventor [46]. investigation [65]. Irrational [95, 93, 96, 101, 102, 100]. irrationnelles [96]. irrationnels [102]. isomorphism [114].

J. [43]. Jacob [42]. January [105]. Johannes [50, 50]. justification [142].

Kurze [51].

Language [119, 74, 75, 146]. languages [131]. Langues [131]. Law [135, 65]. library [137]. life [95]. Linguistic [86, 136, 102]. linguistics [111]. linguistique [102]. locutional [155]. Logic [149, 60, 32, 93, 146]. Logica [149]. Logical [114, 9, 4, 155, 70, 38]. Logique [4, 60]. logiques [9, 70]. Loria [51, 51].

machines [22]. Main [79]. manipulation [21]. Mannoury [105, 106, 103, 108, 104, 110, 109, 112, 107]. Mathematical [9, 71, 25, 27, 70, 63, 75, 73]. Mathematics [132, 10, 74, 89, 51, 153, 49]. Mathematik [49, 74, 51]. mathématique [63]. Mathématiques [132, 9, 70]. mathematischen [73, 75]. May [105]. Meaning [141, 36, 118, 128, 113, 128]. mechanics [11, 77]. Medicine [158]. Memorial [150]. Memoriam [116, 42, 151]. Memoriam/Georges [116]. mental [78, 9, 70]. Mentales [9, 70]. method [108, 64, 58]. Méthode [58, 64]. Methodologisches [109]. methodology [3]. methods [68]. mimic [128]. misinterpretation [136]. model [76]. models [11, 77]. Modern [60, 130, 61, 76, 17, 22]. Moderne [130, 61, 60]. modernes [22]. modernizing [37]. Modes [83, 83]. mot [22]. movements [137].

nature [38]. Nederlandse [149]. negationless [10]. nerve [26]. Neutralité [84]. Neutrality [84]. Nieuwentijt [47, 48]. Non [82, 1, 88]. Non-scientific [1, 88]. Non-verbal [82]. notamment [96, 102]. Note [130, 130].

Obituary [2]. object [40]. objectivité [91]. objectivity [91]. objet [40]. obligations [133]. observables [63]. oddities [155]. Operational [20, 58, 3, 61]. operationalism [62]. Operationalismus [62]. opérationnelle [58]. operationnelles [61]. operations [59, 59]. Ordinary [146]. Ordinary-language [146]. organization [29]. Origin [81]. past [51]. patterns [127]. people [102]. periodicals [7, 122, 139, 157]. peuples [102]. phenomenon [45]. Philosophical [7, 137, 90]. philosophie [91]. Philosophisches [109]. Philosophy [149, 47, 91, 119, 52, 146, 107]. phonetics [99, 100]. Physical [149, 76]. physicist [43]. Physics [110, 130, 79, 61, 129]. physiological [25]. Physique [110, 130, 61]. plausible [153]. Pólya [153]. practice [65]. pratique [65]. precisation [12]. Preface [55]. present [79]. present-day [79]. principles [136]. private [120]. Problem [135, 4]. Problème [135, 4]. problèmes [22]. problems [86, 69, 22, 75]. Problemstellung [75]. process [76, 30]. proof [48]. provoke [102]. provoquent [102]. psychiatry [98]. Psycho [136, 102]. Psycho-linguistic [136, 102]. psycho-linguistique [102]. psycho-sociologie [102]. psycho-sociology [102]. psychology [97].

Quantique [110]. Quantum [110, 11, 77]. Quelques [91, 96, 102]. questionnaires [114]. qu’ils [102]. quotes [51].

rapport [9, 70]. reactions [96, 102, 96, 102]. real [59, 131]. reasoning [153]. réel [59]. Réelles [131]. règles [102]. Reichenbach [2]. reiteration [117]. related [128]. Relation [9, 70]. remarks [91, 51]. remarques [91]. representation [63, 63]. research [86]. resistance [59, 59]. responses [82]. results [99]. Résumé [87]. reviews [41, 121, 156, 53, 13, 8]. role [92, 93, 101, 100]. rule [117]. rules [102].

sans [131]. scarcities [155]. Science [140, 4, 47, 90, 1, 92, 91, 119, 88, 76, 154, 72]. sciences [91, 149]. scientific [1, 88, 19]. Semantic [85, 17, 31]. semantical [118, 120]. semantics [32]. Shannon [125]. Short [51]. sich [51]. sign [128, 75]. signific [17]. Significance [110, 47, 67]. significist [150]. Significs [71, 111, 69, 107]. Signifik [67]. Signifique [110]. signs [81]. simplification [33]. simultaneously [49]. Sinn [128]. slave [96]. Slaves [102]. Slavic [96]. Slavs [102]. slogan [120]. smoking [52]. social [86, 102]. sociales [102]. societies [147]. Society [158]. Sociological [112, 85]. sociologie [83, 102]. Sociology [133, 83, 101, 102, 28]. Some [90, 16, 99, 96, 20, 95, 102, 91]. sound [127]. speech [33, 35, 127, 128]. Sprache [74]. Sprachmimischen [128]. statements [98, 97]. statistical [152]. stimuli [82]. structure [94]. studied [114]. studies [27, 85]. study [80]. subject [128]. subject-related [128]. subjective [91]. subjectives [91]. Summary [87]. syllogistic [143]. Syllogistics [50]. Syllogistik [50]. Symbolic [9, 70]. Symboliques [9, 70]. symbols [1]. synonymity [114]. system [120, 128]. systems [78].

Tautology [5]. teaching [71]. teleological [48]. teleologische [48]. Tendances [60]. tendencies [79]. tessellations [5]. tests [118]. Their [73, 21]. theories [17, 95, 22, 22]. theory [21, 126]. think [51]. thirty [56]. torch [52]. trait [113]. Transformation [33]. Trends [60]. Tribute [104]. tried [51]. Turing [22]. Typen [81]. types [81].

Unbekanntes [49]. unit [72]. Unité [4]. Unity [4]. universals [145]. unknown [49]. Ursprung [81]. usage [36].

Validity [39]. variable [72]. variabler [72]. verbal [82]. Vereniging [149]. Vergangenheit [51]. Verification [98, 97]. versucht [51]. Vertrautes [49]. vocables [102]. volume [115]. Voor [149].

Ways [80]. Wetenschappen [149]. wie [51]. Wijsbegeerte [149]. without [131]. word [22]. work [152]. writing [33]. writings [106].

years [56].

Zeichen [81]. Zeichensprache [75]. Zeichensystem [128]. Zitate [51]. zugleich [49]. Zukunft [51]. Zur [109].

References

Frank:1953:NSS

[1] Philipp G. Frank. Non-scientific symbols in science. Synthese, 9(1):3–9, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

     0964 (electronic). URL http://link.springer.com/article/10.1007/
     BF00567386.

Clay:1953:OHR

[2] J. Clay. Obituary Hans Reichenbach. Synthese, 9(1):10–12, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

     URL http://link.springer.com/article/10.1007/BF00567387.

Somenzi:1953:EOM

[3] Vittorio Somenzi. An exemplification of “operational methodology”. Synthese, 9(1):15–25, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

     article/10.1007/BF00567388.

Ruytinx:1953:DDP

[4] Jacques Ruytinx. Le développement du problème de l’unité de la science dans l’empirisme logique. (French) [The development of the problem of the unity of science in logical empiricism]. Synthese, 9(1):26–41, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

     URL http://link.springer.com/article/10.1007/BF00567389.

Henry:1953:TT

[5] D. P. Henry. Tautology tessellations. Synthese, 9(1):42–47, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00567390.

Nuchelmans:1953:ACC

[6] Gabriel Nuchelmans. The analysis of counterfactual conditionals. Synthese, 9(1):48–63, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567391.

Anonymous:1953:PP

[7] Anonymous. Philosophical periodicals. Synthese, 9(1):64, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00567392;
http://link.springer.com/content/pdf/10.1007/BF00567392.pdf.

Sicard:1953:BR

[8] Emile Sicard. Book reviews. Synthese, 9(1):67–70, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567393.

Piaget:1953:AMR

[9] Jean Piaget. Les activités mentales en rapport avec les expressions symboliques logiques et mathématiques. (French) [Mental activities in relation to logical and mathematical symbolic expressions]. Synthese, 9(1): 73–90, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567394.

Heyting:1953:GFC

[10] A. Heyting. G. F. C. Griss and his negationless intuitionistic mathematics. Synthese, 9(1):91–96, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567395.

Groenewold:1953:QMM

[11] H. J. Groenewold. Quantum mechanics and its models. Synthese, 9(1): 97–103, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567396.

Grimm:1953:ACP

[12] Laura Grimm. On the application of the concept of precisation. Synthese, 9(1):104–120, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567397.

Manders:1953:BR

[13] W. Manders and D. Vuysje. Book reviews. Synthese, 9(1):123–126, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567398.

Anonymous:1953:C

[14] Anonymous. Corrigendum. Synthese, 9(1):128, January 1953-1955. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567399;   http:
//link.springer.com/content/pdf/10.1007/BF00567399.pdf.

Esser:1953:IA

[15] P. H. Esser. Inaugural address. Synthese, 9(1):133–136, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00567400.

vanDantzig:1953:SII

[16] D. van Dantzig. Some informal information on “information”. Synthese, 9(1):137–144, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567401.

Mannoury:1953:SSA

[17] G. Mannoury and D. Vuysje. Semantic and signific aspects of modern theories of communication. Synthese, 9(1):147–156, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567402.

Rapoport:1953:WI

[18] Anatol Rapoport. What is information? Synthese, 9(1):157–173, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567403.

vanSoest:1953:ISI

[19] J. L. van Soest. An introduction to scientific information. Synthese, 9(1): 177–181, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567404.

Mackay:1953:OAS

[20] D. M. Mackay. Operational aspects of some fundamental concepts of human communication. Synthese, 9(1):182–198, January 1953-1955. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567405.

Schouten:1953:CEM

[21] J. F. Schouten. On the concepts of endechy and manipulation and their application to information theory. Synthese, 9(1):199–204, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00567406.

Tamari:1953:CTM

[22] Dov Tamari. Une contribution aux théories modernes de communication: machines de Turing et problèmes de mot. (French) [A contribution to modern communication theories: Turing machines and word problems]. Synthese, 9(1):205–227, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567407.

Fortuyn:1953:I

[23] J. Droogleever Fortuyn. Introduction. Synthese, 9(1):231–232, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com//content/pdf/10.1007/
BF00567408.pdf;   http://link.springer.com/article/10.1007/BF00567408.

McCulloch:1953:IH

[24] Warren S. McCulloch. Information in the head. Synthese, 9(1):233–247, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567409.

Karreman:1953:MBP

[25] George Karreman. The mathematical biology of physiological excitation. Synthese, 9(1):248–251, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567410.

deVries:1953:ICH

[26] Hl. de Vries. The information capacity of the human ear and of the cochlear nerve. Synthese, 9(1):252–254, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00567411.

Karreman:1953:RMB

[27] George Karreman. Recent mathematical–biological studies on communication. Synthese, 9(1):255–261, January 1953-1955. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567412.

vanSoest:1953:CIS

[28] J. L. van Soest. A contribution of information to sociology. Synthese, 9(1): 265–273, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567413.

Poortman:1953:FO

[29] B. Poortman. Foundations of organization. Synthese, 9(1):274–288, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567414.

Suri:1953:CP

[30] Surindar Suri. The communicative process. Synthese, 9(1):289–295, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567415.

Bar-Hillel:1953:ICS

[31] Yehoshua Bar-Hillel. Information and content: A semantic analysis. Synthese, 9(1):299–305, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567416.

Reiser:1953:LCS

[32] Oliver L. Reiser. Logic, cybernetics, and semantics. Synthese, 9(1):306– 318, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567417.

DeWitte:1953:TSW

[33] A. J. J. De Witte. Transformation of speech into writing and simplification. Synthese, 9(1):321–325, January 1953-1955. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567418.

Vries:1953:CI

[34] M. De Vries. Concealment of information. Synthese, 9(1):326–336, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567419.

Esser:1953:FS

[35] P. H. Esser. Final speech. Synthese, 9(1):339, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (elec-

tronic). URL http://link.springer.com//content/pdf/10.1007/
BF00567420.pdf;   http://link.springer.com/article/10.1007/BF00567420.

Gullvaag:1953:CMA

[36] Ingemund Gullvåg. Criteria of meaning and analysis of usage. Synthese, 9(1):341–361, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567421.

Crockett:1953:DBM

[37] Campbell Crockett. The distinction between modernizing and dissolving aesthetics. Synthese, 9(1):362–371, January 1953-1955. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00567422.

Skolem:1953:LNA

[38] Th. Skolem. The logical nature of arithmetic. Synthese, 9(1):375–384, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567423.

Vermeulen:1953:VH

[39] R. Vermeulen. Validity of hypotheses. Synthese, 9(1):385–394, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567424.

Berghuys:1953:LGF

[40] J. J. W. Berghuys. L’objet de la géometrie. (French) [The object of geometry]. Synthese, 9(1):395–407, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567425.

Beth:1953:BR

[41] E. W. Beth, Esser, H. J. Groenewold, and Leo Apostel. Book reviews. Synthese, 9(1):411–416, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567426.

Anonymous:1953:MJC

[42] Anonymous. In memoriam Jacob Clay. Synthese, 9(1):423–427, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567427.

Jongen:1953:PJC

[43] H. F. Jongen. The physicist J. Clay. Synthese, 9(1):428–432, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567428.

Frank:1953:E

[44] Philipp G. Frank. Einstein. Synthese, 9(1):435–437, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00567429.

Mannoury:1953:CPA

[45] G. Mannoury. The cultural phenomenon Albert Einstein. Synthese, 9(1): 438–441, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567430.

Fokker:1953:AEI

[46] A. D. Fokker. Albert Einstein, inventor of chronogeometry. Synthese, 9(1): 442–445, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567431.

Beth:1953:NSP

[47] E. W. Beth. Nieuwentijt’s significance for the philosophy of science. Synthese, 9(1):447–453, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567432.

Freudenthal:1953:NTG

[48] Hans Freudenthal. Nieuwentijt und der teleologische Gottesbeweis. (German) [Nieuwentijt and the teleological proof of God]. Synthese, 9(1): 454–464, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567433.

Bernays:1953:MZV

[49] Paul Bernays. Die Mathematik als ein zugleich Vertrautes und Unbekanntes. (German) [Mathematics as a simultaneously familiar and unknown]. Synthese, 9(1):465–471, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00567434.

Durr:1953:SJH

[50] Karl Dürr. Die Syllogistik des Johannes Hospinianus (1515–1575). (German) [The syllogistics of Johannes Hospinianus (1515–1575)]. Synthese, 9(1):472–484, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00567435.

Vollgraff:1953:KBZ

[51] J. A. Vollgraff. Kurze Bemerkungen, und Zitate, über Vergangenheit und Zukunft der Mathematik, so wie Gino Loria (1862–1954) versucht hat sich Dieselben zu Denken. (German) [Short remarks, and quotes, on the past and future of mathematics, as Gino Loria (1862–1954) has tried to think of Diesel]. Synthese, 9(1):485–491, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00567436.

Plooij:1953:TPS

[52] E. B. Plooij. The torch of philosophy, a smoking flax. Synthese, 9(1): 492–498, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00567437.

Kaiser:1953:BR

[53] L. Kaiser and D. Vuysje. Book reviews. Synthese, 9(1):501–503, January 1953-1955. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00567438.

Anonymous:1956:A

[54] Anonymous. Acknowledgments. Synthese, 10(1):4, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484631.

Anonymous:1956:PI

[55] Anonymous. Preface and introduction. Synthese, 10(1):5–9, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484632.

Mannoury:1956:ATY

[56] G. Mannoury. After thirty years. Synthese, 10(1):13–14, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484633.

Frank:1956:IA

[57] Philipp Frank. Introductory address. Synthese, 10(1):15–19, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484634.

Piaget:1956:MAM

[58] Jean Piaget. Méthode axiomatique et méthode opérationnelle. (French) [Axiomatic method and operational method]. Synthese, 10(1):23–43, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484635.

Inhelder:1956:GIH

[59] Bärbel Inhelder. La genèse de l’idée de hasard et la résistance du réel aux opérations. (French) [The genesis of the idea of chance and the resistance of the real to operations]. Synthese, 10(1):44–50, January 1956-1958. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484636.

Fevrier:1956:TCL

[60] P. Fevrier. Tendances constructives en logique moderne. (French) [Constructive trends in modern logic]. Synthese, 10(1):52–58, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484637.

Destouches:1956:DOP

[61] Jean-Louis Destouches. Descriptions operationnelles en physique moderne. (French) [Operational descriptions in modern physics]. Synthese, 10(1):59–64, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484638.

Szekely:1956:BOG

[62] D. L. Székely. Begriffsbildung und Operationalismus. (German) [Conceptualisation and operationalism]. Synthese, 10(1):65–70, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484639.

Servien:1956:RMO

[63] Pius Servien. La représentation mathématique des observables. (French) [The mathematical representation of observables]. Synthese, 10(1):71–75, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484640.

Bourgin:1956:MH

[64] Georges Bourgin. La méthode historique. (French) [The historical method]. Synthese, 10(1):79–83, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484641.

Houin:1956:EAD

[65] Roger Houin. Une enquête sur l’application du droit dans la pratique. (French) [An investigation into the application of law in practice]. Synthese, 10(1):84–90, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484642.

deWitte:1956:TNG

[66] A. J. J. de Witte. Towards a new grammar. Synthese, 10(1):91–101, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484643.

Denk:1956:SHG

[67] F. Denk. Signifik und Heuristik. (German) [Significance and Heuristics]. Synthese, 10(1):102–107, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484644.

Clay:1956:CM

[68] J. Clay. Combination of methods. Synthese, 10(1):111–119, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484645.

Mannoury:1956:DPS

[69] G. Mannoury. Didactic problems in significs. Synthese, 10(1):123–126, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484646.

Piaget:1956:AMR

[70] Jean Piaget. Les activités mentales en rapport avec les expressions symboliques logiques et mathématiques. (French) [Mental activities in relation to logical and mathematical symbolic expressions]. Synthese, 10(1):127– 145, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484647.

Geursen:1956:SMT

[71] S. J. Geursen. Significs and mathematical teaching. Synthese, 10(1):146– 147, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484648.

Szekely:1956:FED

[72] David L. Székely. Formalisierung der Einheitswissenschaft mit Dimensionsbegriffen von variabler Exaktheit. (German) [Formalization of unit science with dimensions of variable accuracy]. Synthese, 10(1):148–170, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484649.

Denk:1956:MBI

[73] F. Denk. Die mathematischen Begriffe und ihr Ausdruck. (German) [Mathematical concepts and their expression]. Synthese, 10(1):173–180, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484650.

Lorenzen:1956:MSG

[74] Paul Lorenzen. Ist Mathematik eine Sprache?. (German) [Is mathematics a language?]. Synthese, 10(1):181–186, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484651.

Walter:1956:AMZ

[75] Emil J. Walter. Die Ausbildung der mathematischen Zeichensprache und ihr Einfluss auf die Entwicklung der mathematischen Problemstellung. (German) [The formation of mathematical sign language and its influence on the development of mathematical problems]. Synthese, 10(1): 187–189, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484652.

Holton:1956:HMG

[76] Gerald Holton. A heuristic model for the growth process of modern physical science. Synthese, 10(1):190–202, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484653.

Groenewold:1956:QMM

[77] H. J. Groenewold. Quantum mechanics and its models. Synthese, 10(1): 203–209, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484654.

Kazemier:1956:FSM

[78] B. H. Kazemier. Formal systems and mental activity. Synthese, 10(1): 210–219, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484655.

Clay:1956:MTP

[79] J. Clay. Main tendencies in present-day physics. Synthese, 10(1):220– 225, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484656.

Witte:1956:WID

[80] A. J. J. De Witte. Ways of interpretation. A didactic study. Synthese, 10(1):229–237, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484657.

Dohmann:1956:UET

[81] Karl Döhmann. Ursprung und Entwicklung, Typen und Funktionen der Zeichen. (German) [Origin and development, types and functions of signs]. Synthese, 10(1):238–245, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484658.

Esser:1956:NVR

[82] P. H. Esser. Non-verbal responses to verbal stimuli. Synthese, 10(1): 246–258, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484659.

Leenhardt:1956:MES

[83] Maurice Leenhardt. Modes d’expression en sociologie et en ethnologie. (French) [Modes of expression in sociology and ethnology]. Synthese, 10(1):259–264, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484660.

Bourgin:1956:N

[84] Georges Bourgin. Neutralité. (French) [Neutrality]. Synthese, 10(1):265– 269, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484661.

Leyton:1956:SAS

[85] A. C. Leyton. Semantic aspects of sociological studies. Synthese, 10(1): 270–278, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484662.

Cervin:1956:LPR

[86] Vladimir Cervin. Linguistic problems in recent social research. Synthese, 10(1):279–281, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484663.

Perelman:1956:RIF

[87] Ch. Perelman. Résumé des interventions. (French) [Summary of interventions]. Synthese, 10(1):285–289, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484664.

Groenewold:1956:NSE

[88] H. J. Groenewold. Non-scientific elements in the development of science. Synthese, 10(1):293–311, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484665.

Mannoury:1956:IBF

[89] G. Mannoury. The ideological background of the foundation of mathematics. Synthese, 10(1):315–317, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484666.

Bridgman:1956:SPA

[90] P. W. Bridgman. Some philosophical aspects of science. Synthese, 10(1): 318–326, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484667.

Gagnebin:1956:QRS

[91] S. Gagnebin. Quelques remarques subjectives sur l’objectivité en sciences et en philosophie. (French) [Some subjective remarks on objectivity in science and philosophy]. Synthese, 10(1):327–334, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484668.

Frank:1956:RAI

[92] Philipp Frank. The role of authority in the interpretation of science. Synthese, 10(1):335–338, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484669.

Jorgensen:1956:RID

[93] Jørgen Jørgensen. The role of the irrational in the development of logic. Synthese, 10(1):341–348, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484670.

Apostel:1956:FSA

[94] L. Apostel. The formal structure of action. Synthese, 10(1):349–356, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484671.

Raven:1956:IES

[95] Chr. P. Raven. Irrational elements in some theories of life. Synthese, 10(1): 359–363, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484672.

Lajberich:1956:QRE

[96] C.-M. Lajberich. De quelques réactions émotionnelles irrationnelles chez l’enfant et notamment chez l’enfant slave. (French) [On some irrational emotional reactions in the child and especially in the Slavic child]. Synthese, 10(1):364–368, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484673.

Vuysje:1956:VSP

[97] D. Vuysje. A. “Verification” of statements in psychology. Synthese, 10(1): 369–372, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484674.

Esser:1956:BVS

[98] P. H. Esser. B. “Verification” of statements in psychiatry. Synthese, 10(1): 373–377, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484675.

Kaiser:1956:ASR

[99] L. Kaiser. Aspects of some results of experimental phonetics. Synthese, 10(1):381–384, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484676.

deWitte:1956:RID

[100] A. J. J. de Witte. The role of the irrational in the development of phonetics. Synthese, 10(1):385–386, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484677.

Leyton:1956:RIS

[101] Alvin C. Leyton. The role of the irrational in sociology. Synthese, 10(1): 389–398, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484678.

Sicard:1956:QRI

[102] Emile Sicard. De quelques règles d’interprétation des éléments irrationnels de la psycho-sociologie des peuples par l’analyse psycho-linguistique des vocables et des réactions sociales qu’ils provoquent, notamment chez les slaves. (French) [Ib some rules of interpretation of the irrational elements of the psycho-sociology of people through the psycho-linguistic analysis of the vocables and of the social reactions they provoke, especially among the Slavs]. Synthese, 10(1):399–407, January 1956-1958. CODEN SYN-TAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484679.

Anonymous:1956:GM

[103] Anonymous. Gerrit Mannoury. Synthese, 10(1):409–413, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484680.

Bourgin:1956:HGM

[104] Georges Bourgin. Hommage à Gerrit Mannoury. (French) [Tribute to Gerrit Mannoury]. Synthese, 10(1):414–415, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484681.

Anonymous:1956:GMM

[105] Anonymous. Gerrit Mannoury May 17, 1867–January 30, 1956. Synthese, 10(1):416–417, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484682.

Anonymous:1956:BWG

[106] Anonymous. Bibliography of the writings of Gerrit Mannoury. Synthese, 10(1):418–422, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484683.

vanDantzig:1956:MIP

[107] D. van Dantzig. Mannoury’s impact on philosophy and significs. Synthese, 10(1):423–431, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484684.

Beth:1956:MM

[108] E. W. Beth. On Mannoury’s method. Synthese, 10(1):432–439, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484685.

Heyting:1956:MMP

[109] A. Heyting. Mannoury’s “Methodologisches Und Philosophisches Zur Elementarmathematik”. Synthese, 10(1):440–444, January 1956-1958. CO-DEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484686.

Destouches:1956:SMC

[110] Jean Louis Destouches. La signifique de Mannoury comparée à la physique quantique et à la cybernétique. (French) [The significance of Mannoury compared to quantum physics and cybernetics]. Synthese, 10(1):445–447, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484687.

deWitte:1956:SL

[111] A. J. J. de Witte. Significs and linguistics. Synthese, 10(1):448–453, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484688.

Meertens:1956:SCG

[112] P. J. Meertens. Sociological conceptions of G. Mannoury. Synthese, 10(1): 454–460, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484689.

deGroot:1956:MTC

[113] A. D. de Groot. The meaning of trait concepts. Synthese, 10(1):461– 470, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484690.

Naess:1956:LEI

[114] Arne Naess. Logical equivalence, intentional isomorphism and synonymity as studied by questionnaires. Synthese, 10(1):471–479, January 1956-1958. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484691.

Anonymous:1959:INV

[115] Anonymous. Introduction to the new volume. Synthese, 11(1):3–4, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486027;
http://link.springer.com/content/pdf/10.1007/BF00486027.pdf.

Levy-Bruhl:1959:MGB

[116] Henri Levy-Bruhl. In memoriam/Georges Bourgin. Synthese, 11(1): 5–6, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486028.

Dodd:1959:RR

[117] Stuart C. Dodd. The reiteration rule. Synthese, 11(1):7–32, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486029.

Hiorth:1959:DMS

[118] Finngeir Hiorth. Distances of meaning and semantical tests. Synthese, 11(1):33–62, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486030.

George:1959:LPE

[119] F. H. George. Language, philosophy and empirical science. Synthese, 11(1):63–71, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486031.

Tennessen:1959:SPE

[120] Herman Tennessen. ‘The system of private enterprise’, an empiriosemantical analysis of a slogan. Synthese, 11(1):72–83, March 1959. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486032.

Beth:1959:BR

[121] E. W. Beth, Edward W. Barankin, A. Cohen, and J. L. van Soest. Book reviews. Synthese, 11(1):84–94, March 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486033.

Anonymous:1959:RPa

[122] Anonymous. Recent periodicals. Synthese, 11(1):95–96, March 1959. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486034.

Anonymous:1959:I

[123] Anonymous. Introduction. Synthese, 11(2):101–103, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485569.

vanSoest:1959:II

[124] J. L. van Soest. Information and interpretation. Synthese, 11(2):104– 111, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485570.

Loeb:1959:CAS

[125] Julien Loeb. Le calcul de l’ambiguité de Shannon en code binaire. (French) [Calculation of the Shannon ambiguity in binary code]. Synthese, 11(2): 112–118, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485571.

Stumpers:1959:ICT

[126] F. L. H. M. Stumpers. Interpretation and communication theory. Synthese, 11(2):119–126, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485572.

Kaiser:1959:ISP

[127] Louise Kaiser. Informative sound patterns in speech. Synthese, 11(2): 127–136, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485573.

Trojan:1959:BGS

[128] Felix Trojan. Bedeutung, Gegenstandsbeziehing und Sinn im Sprachmimischen Zeichensystem. (German) [Meaning, subject-related and meaning in the speech–mimic–sign system]. Synthese, 11(2):137–149, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485574.

Somenzi:1959:HIA

[129] Vittorio Somenzi. Human interpretation and automatic interpretation in physics. Synthese, 11(2):150–152, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485575.

Mercier:1959:NIC

[130] André Mercier. Note sur l’interprétation de la causalité en physique moderne. (French) [Note on the interpretation of causality in modern physics]. Synthese, 11(2):153–159, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485576.

Mandelbrot:1959:ISI

[131] Benoît Mandelbrot. Information sans interprétation dans la déscription des langues réelles. (French) [Information without interpretation in the description of real languages]. Synthese, 11(2):160–166, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic).  URL http://link.springer.com/article/10.1007/
BF00485577;   http://www.jstor.org/stable/20114289;   http:/
/www.springerlink.com/content/r10332532461x0r7/.

Tamari:1959:IMF

[132] Dov Tamari. Interprétations en mathématiques. (French) [Interpretations in mathematics]. Synthese, 11(2):167–176, June 1959. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485578.

Leyton:1959:SOI

[133] Alvin C. Leyton. Sociology and obligations of interpretation. Synthese, 11(2):177–196, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485579.

Crockett:1959:II

[134] Campbell Crockett. On interpretations of ‘interpretation’. Synthese, 11 (2):197–214, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485580.

Houin:1959:PID

[135] Roger Houin. Le problème d’interprétation dans le droit. (French) [The problem of interpretation in law]. Synthese, 11(2):215–219, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485581.

DeWitte:1959:PLP

[136] A. J. J. De Witte. Psycho-linguistic principles and backgrounds of misinterpretation. Synthese, 11(2):220–222, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485582.

Anonymous:1959:LPM

[137] Anonymous. The library of philosophical movements. Synthese, 11(2): 223, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485583.

Anonymous:1959:Ca

[138] Anonymous. Congresses. Synthese, 11(2):223, June 1959. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485584.

Anonymous:1959:RPb

[139] Anonymous. Recent periodicals. Synthese, 11(2):224–226, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485585.

Beth:1959:SC

[140] E. W. Beth. Science and classification. Synthese, 11(3):231–244, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486414.

George:1959:MB

[141] F. H. George. Meaning and behaviour. Synthese, 11(3):245–258, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486415.

Margolis:1959:DJI

[142] Joseph Margolis. The demand for a justification of induction. Synthese, 11 (3):259–264, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486416.

Prior:1959:FS

[143] A. N. Prior. Formalised syllogistic. Synthese, 11(3):265–273, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486417.

Tennessen:1959:EI

[144] Herman Tennessen. Evidence and illustration. Synthese, 11(3):274–276, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486418.

Weissmann:1959:EU

[145] H. A. Weissmann. The essence of universals. Synthese, 11(3):277–293, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486419.

Xenakis:1959:OLP

[146] Jason Xenakis. Ordinary-language philosophy: Language, logic and philosophy. Synthese, 11(3):294–306, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00486420.

Anonymous:1959:NS

[147] Anonymous. New societies. Synthese, 11(3):307, September 1959. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486421;   http:
//link.springer.com/content/pdf/10.1007/BF00486421.pdf.

Anonymous:1959:Cb

[148] Anonymous. Congresses. Synthese, 11(3):308, September 1959. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486422;   http:
//link.springer.com/content/pdf/10.1007/BF00486422.pdf.

Anonymous:1959:NVV

[149] Anonymous. Nederlandse Vereniging Voor Logica en Wijsbegeerte der Exacte Wetenschappen. (Dutch) [Dutch Association for Logic and Philosophy of Physical Sciences]. Synthese, 11(3):309–310, September 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486423.

Anonymous:1959:MDV

[150] Anonymous. Memorial to David van Dantzig: van Dantzig as a significist. Synthese, 11(4):319–328, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486193.

Koksma:1959:MDV

[151] J. F. Koksma. In memoriam David van Dantzig. Synthese, 11(4):329– 334, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486194.

Hemelrijk:1959:DVD

[152] J. Hemelrijk. David van Dantzig’s statistical work. Synthese, 11(4):335– 351, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486195.

vanDantzig:1959:GPM

[153] D. van Dantzig. G. Pólya mathematics and plausible reasoning. Synthese, 11(4):353–358, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486196.

Smart:1959:CBE

[154] J. J. C. Smart. Can biology be an exact science? Synthese, 11(4):359– 368, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486197.

Tennessen:1959:LOL

[155] Herman Tennessen. Logical oddities and locutional scarcities. Synthese, 11 (4):369–388, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486198.

Groenewold:1959:BR

[156] H. J. Groenewold, P. H. Esser, and N. de Raadt. Book reviews. Synthese, 11(4):389–400, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486199.

Anonymous:1959:RPc

[157] Anonymous. Recent periodicals. Synthese, 11(4):401–404, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486200.

Anonymous:1959:ISC

[158] Anonymous. International society of cybernetic medicine. Synthese, 11 (4):405, December 1959. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486201.